1/1

Lord George Gordon Byron ,

My soul is dark

My soul is dark – Oh! quickly string The harp I yet can brook to hear; And let thy gentle fingers fling Its melting murmurs o’er mine ear.

If in this heart a hope be dear, That sound shall charm it forth again: If in these eyes there lurk a tear, ’Twill flow, and cease to burn my brain.

But bid the strain be wild and deep, Nor let thy notes of joy be first; I tell thee, minstrel, I must weep, Or else this heavy heart will burst;

For it hath been by sorrow nursed, And ached in sleepless silence long; And now ’tis doom’d to know the worst, And break at once – or yield to song.

Verkinto de tiu ĉi Angla poemo estas Lord George Gordon Byron (∗1788-01-22 – †1824-04-18).

Arg-1018-2060 (2014-05-20 20:08:48)

Pri la poeto vidu la vikipediejon http: // eo. wikipedia. org/ wiki/ George_Byron . Jen ankaŭ la retejo de la Bajron-societo, Byron-society: http: // www. thebyronsociety. com/ contact. html .

http://www.poezio.net/version?poem-id=1018&version-id=,2060